1. When 15.4 g of aluminum (26.98 g/mol) reacts with 14.8 g of oxygen gas (32.00 g/mol), what is the maximum mass (in grams) of aluminum oxide (101.96 g/mol) that can be produced? 4 Al(s) + 3 O2(g) → 2 Al2O3(s)

2. If 14.1 moles of Cu and 48.7 moles of HNO3 are allowed to react, how many moles of excess reactant will remain if the reaction goes to completion? Cu + 8 HNO3 → 3Cu(NO3)2 + 2NO + 4H2O

3. Calculate the number of moles of excess reactant that will be left-over when 50.0 g of KI react with 50.0 g of Br2: [ 2KI + Br2   ->   2KBr + I2 ]
Answer #1

1. From reaction

15.4 g Al *(1 mol Al/ 26.98 g)*(3 mol O2 / 4 mol Al)*(32 g/ 1 mol) = 13.70 g O2

Thus, Oxygen is a limiting reagent as amount of O2 present is less

Now, Maximum mass of Al2O3 present;

14.8 g O2 *( 1 mol/ 32 g )*(2 mol Al2O3 / 3 mol O2)*(101.96 g/ 1 mol) = 31.4 g Al2O3

2. From reaction

14.1 mol Cu * ( 8 mol HNO3 / 1 mol Cu) = 112.8 mol HNO3

Thus, Cu in excess

48.7 mol HNO3 * ( 1 mol Cu / 8 mol HNO3) = 6.09 mol Cu

Excess mol of Cu = 14.8 - 6.09 = 8.71 mol of Cu

3. From reaction

50.0 g KI * ( 1 mol KI / 166 g)*( 1 mol Br2 / 2 mol KI) * (159.8 g Br2 / 1 mol) = 24.1 g Br2

Thus, Br2 is in excess and excess amount of Br2 = 50.0 - 24.1 = 25.9 g / 159.8g/mol = 0.16 mol